### v1

# Tractable Algorithms for Changepoint Detection in Player Performance Metrics

Abstract
We present tractable methods for detecting changes in player performance metrics and apply these methods to Major League Baseball (MLB) batting and pitching data from the 2023 and 2024 seasons. First, we derive principled benchmarks for when performance metrics can be considered statistically reliable, assuming no underlying change, using distributional assumptions and standard concentration inequalities. We then propose a changepoint detection algorithm that combines a likelihood-based approach with split-sample inference to control false positives, using either nonparametric tests or tests appropriate to the underlying data distribution. These tests incorporate a shift parameter, allowing users to specify the minimum magnitude of change to detect. We demonstrate the utility of this approach across several baseball applications: detecting changes in batter plate discipline metrics (e.g., chase and whiff rate), identifying velocity changes in pitcher fastballs, and validating velocity changepoints against a curated ground-truth dataset of pitchers who transitioned from relief to starting roles. Our method flags meaningful changes in 91% of these ‘ground-truth’ cases and reveals that, for some metrics, more than 60% of detected changes occur in-season. While developed for baseball, the proposed framework is broadly applicable to any setting involving monitoring of individual performance over time.
1 Introduction
Aaron Judge, an outfielder for the New York Yankees, led Major League Baseball (MLB) in on-base percentage (OBP) during the 2024 season, finishing with an impressive OBP of .458 compared to the league average of .312. Early in the season, however, his OBP after his first 24 plate appearances (PA) was just .208, substantially lower than his end-of-season OPB. Intuitively, this discrepancy highlights the well-known issue that performance metrics based on small sample sizes can vary substantially, and that this variability typically decreases as sample size increases. This example naturally raises a fundamental question relevant not just in sports analytics but in any field that tracks performance over time: how can we efficiently detect meaningful changes in a metric, or determine when apparent changes are simply due to small sample sizes?
Across sports, sample size in relation to metric stabilization and changepoints is a widely discussed topic with researchers often asking how large a sample is needed to be reasonably confident in the corresponding results (Carleton 2012, Foley 2024). This concern about small sample variability is foundational in statistics, with the James-Stein estimator providing a classic example in baseball: even seemingly straightforward (batting) averages can be misleadingly variable when based on limited data (Efron & Morris 1977). In this paper, we formalize the concept of metric stabilization, proposing methods for determining when a metric becomes sufficiently stable in baseball, and introduce algorithms for detecting changes in player performance metrics over time.
We note that calculations to determine metric stabilization points should not be entirely based on player data (especially spanning multiple seasons), as players often change over the course of the season, as well as across seasons. In fact, our changepoint analysis (see Sections 4-6) indicates that for some statistics more than 60% of detected changes occur outside of the first or last month of the season (i.e., occuring in May through August).
We present efficient algorithms for detecting changepoints in both binary and continuous performance metrics at the individual player level. While changepoint detection methods have been used in prior sports analytics research, they have primarily focused on identifying changes in aggregated statistics for the average player, team, or league across seasons (Chiou et al. 2016, Whalen et al. 2024). In contrast, our focus is on identifying changes in individual performance across time, which poses a unique computational challenge due to the large number of players to monitor in real time.
Our approach draws on ideas from the changepoint detection literature (Hinkley & Hinkley 1970, Sen & Srivastava 1975, Killick & Eckley 2014, Matteson & James 2014, Shang et al. 2017, Jewell et al. 2022) but integrates them in a novel framework specifically designed for this application. In particular, we incorporate split-sample inference with hypothesis tests that include a shift parameter, giving users direct control over both the Type I error rate and the magnitude of changes considered meaningful. This addresses a key limitation in existing R packages (Lindeløv 2024), which typically assume normally distributed data and omit split-sample inference, and therefore tend to overflag changepoints and perform poorly for binary outcomes like plate discipline metrics. We provide open-source R functions that handle both continuous and binary data efficiently, using a likelihood ratio statistic maximized across all possible changepoints for a given player. We validate our method on a constructed ground-truth dataset of pitchers who changed roles from reliever to starter during the 2023 or 2024 MLB seasons, and illustrate its broader utility by detecting changes in fastball velocity and plate discipline metrics such as zone swing percentage, whiff percentage, and chase rate.
Although this paper focuses on applications to MLB, the algorithms and software we present are broadly applicable to any domain where detecting changes in individual performance metrics is of interest, such as in manufacturing, where changepoint detection has been a longstanding area of research (Shang et al. 2017).
The paper is organized as follows. Section 2 describes the MLB datasets. In Section 3, we formalize methods for benchmarking metric stability under the assumption of no changepoints and present stabilization points for several common MLB batting statistics. Section 4 introduces our changepoint detection algorithms. We apply these algorithms to batter plate discipline metrics in Section 5, and to pitcher fastball velocity in Section 6. Ground-truth validation is presented in Section 7. Section 8 concludes with a discussion of limitations and directions for future research.
2 MLB Data
We consider batter and pitcher data at the plate appearance- and pitch-level from the 2023 and 2024 MLB seasons. All regular season games were included; postseason games and spring training data were excluded.
For the stabilization analysis in Section 3, plate appearance data was obtained from FanGraphs. A plate appearance (PA) refers to a batter’s completed turn at bat, while an at-bat (AB) is a plate appearance that excludes outcomes that do not count towards batting average, i.e., walks, hit-by-pitches, and sacrifices. We focus on the following batter performance metrics (aggregated across different numbers of PAs across a season):
  • Batting average (BA): hits divided by at-bats
  • On-base percentage (OBP): the percentage of a batter’s plate appearances that result in the batter reaching base
  • Walk rate (BB%): the percentage of a batter’s plate appearances that result in a walk
  • Strikeout rate (K%): the percentage of a batter’s plate appearances that result in a strikeout
  • Hit rate (H%): the percentage of a batter’s plate appearances that result in a hit
  • Homerun rate (HR%): the percentage of a batter’s plate appearances that result in a homerun
For the changepoint analysis in Sections 5-7, we obtained pitch-level data from MLB Statcast, accessed via Baseball Savant. For pitchers, we analyze fastball velocity at the pitch level, treating four-seam fastballs and sinkers as distinct pitch types. For batters, we analyze plate discipline metrics, which evaluate a player’s ability to be selective with their swings, i.e., to swing at pitches they can hit well while taking pitches they cannot. For example, some plate discipline metrics evaluate a player’s ability to swing at pitches within the strike zone and refrain from swinging at pitches outside of the strike zone. The strike zone is defined as the area over home plate from approximately the batter’s knees to the midpoint of their torso, as measured by Statcast. In this paper, we focus on the following batter plate discipline metrics:
  • Out-of-zone swing: whether the batter swings at a pitch located outside the strike zone. Aggregating this metric across pitches (outside the strike zone) yields the out-of-zone swing percentage (O-Swing%), also known as chase rate.
  • Swinging strike: whether the batter fails to make contact with a pitch that they swing at. Aggregating this metric across all pitches a batter swings at yields the swinging strike rate (SwStr%), also known as whiff rate.
3 Benchmarking Metric Stability Under No Change
In sports, the idea of metric “stabilization” is often discussed informally, with many practitioners deriving general thresholds from retrospective player data (Carleton 2012, Foley 2024). However, these heuristics can be misleading, as they typically assume that all observed variation reflects sampling noise, even when a player’s underlying skill may have changed. In this section, we formalize the concept of stabilization by deriving principled thresholds for when a player’s observed performance metric is likely to fall within a specified range of their true value, under the assumption that no change has occurred.
While the no change assumption can be examined using changepoint detection methods introduced later in the paper, the stabilization thresholds derived here offer a complementary diagnostic: they characterize the amount of variation one should expect due to sampling alone. This benchmark allows analysts to better interpret fluctuations, identifying when a change is statistically plausible, and also recognizing when surprising values may simply reflect limited sample size. In later sections, we characterize how frequently player performance metrics do exhibit meaningful changes and how often these changes occur in-season versus during the off season.
3.1 Metric Stabilization
Suppose a player has true underlying value $\theta$ for some metric which we estimate at time $t$ as $T_{t}$. We say that $T_{t}$ has stabilized when it becomes sufficiently close (i.e., within some threshold $S$) to $\theta$. Furthermore, we define the stabilization point as follows:
Definition 3.1 (Stabilization point)
Let a statistic $T_{t}$ be an estimate for $\theta$ at time $t$. We define the stabilization point for $T_{t}$ as the smallest value of $t$ such that the mean-squared error is bounded by $S^{2}$, i.e., $\mathbb{E}[(T_{t}-\theta)^{2}]\leq S^{2}$.
In other words, we are looking for the time point where on average we expect our estimate to be sufficiently close (i.e., within $S$) to our true parameter value.
Because player performance can change over the course of the season (see Sections 5 and 6) we want to define stabilization points for metrics without using player data when we are unsure whether or not a change occurs. Therefore, we will define stabilization points using the distributional properties of the metrics. We also discuss how to construct sequentially valid confidence intervals for metrics throughout the course of the season.
3.2 Bounded Metrics: Rate Statistics
Many player performance metrics in baseball are based on binary outcomes, e.g., on-base percentage (OBP), strikeout rate (K%), and homerun rate (HR%). These metrics can be modeled as averages of independent Bernoulli random variables. In this section, we derive stabilization benchmarks for such metrics by computing their mean squared error and constructing confidence intervals using classical concentration inequalities. We then apply this framework to several standard batting statistics and provide benchmark stabilization points that reflect typical levels of inter-player variability.
3.2.1 Stabilization Point
Consider a player skill, such as home run rate, with true underlying probability $p$ of success. Each event (e.g., plate appearance, at-bat) follows a Bernoulli distribution: $X_{i}\sim\text{Bernoulli}(p).$
The observed rate over $n$ events is given by the sample mean $\hat{p}=\sum_{i}X_{i}/n$ which has variance $Var(\hat{p})=\frac{p(1-p)}{n}$. Since $\hat{p}$ is an unbiased estimate of $p$, the mean-squared error (MSE) reduces to the variance:
$\mathbb{E}[(\frac{\sum_{i}X_{i}}{n}-p)^{2}]=Var(\frac{\sum_{i}X_{i}}{n})=\frac{p(1-p)}{n}$
To find the stabilization point, according to Definition 3.1 we find the smallest $t$ such that
$\frac{p(1-p)}{t}\leq S^{2}\implies t\geq\frac{p(1-p)}{S^{2}}.$
The number of observations required for stabilization depends not only on the precision threshold $S$, but also on the underlying skill level $p$. Stabilization occurs more quickly when $p$ is near 0 or 1, and is slowest when $p=0.5$, where the variance is maximized.
A natural way to set $S$ is as the latent standard deviation of true talent across players, i.e., the true spread of $p$ in the population. The observed variance of a metric across players, $\sigma^{2}_{\text{obs}}$, combines two sources of variation: 1) the true variance of $p$ across players, $\sigma^{2}_{\text{latent}}$, and 2) the variance due to finite-sample noise for each player’s estimate, $\sigma^{2}_{\text{samp}}$. These relate by
$\sigma^{2}_{\text{obs}}=\sigma^{2}_{\text{latent}}+\sigma^{2}_{\text{samp}}$
Therefore, we can estimate S to be one latent standard deviation:
$S=\sqrt{\sigma^{2}_{\text{obs}}-\sigma^{2}_{\text{samp}}}$
Where we define $\sigma_{\text{obs}}$ as the observed standard deviation of $\hat{p}$ across all players in the population and $\sigma_{\text{samp}}$ as the average sampling standard deviation expected due to finite $n$. Defining $n_{\text{stable}}$ as the number of observations required for the root mean squared error (RMSE) of $\hat{p}$ to fall within one latent standard deviation $\sigma_{\text{latent}}$, we have:
$n_{stable}=\frac{\hat{p}(1-\hat{p})}{\hat{\sigma}_{\text{latent}}^{2}}$
Table 1 gives the stabilization values $n_{stable}$ for various batter statistics.
Table 1: Stabilization point $n_{stable}$ for batter statistics: batting average (BA), on-base percentage (OBP), walk rate (BB%), strikeout rate (K%), hit rate (H%), and homerun rate (HR%). For each metric, the average season-ending value across players with more than 3162=4863*162=486 plate appearances is given, as well as the corresponding standard deviation values.
Metric Average $\hat{p}$ $\sigma_{obs}$ $\sigma_{samp}$ $\sigma_{latent}$ $n_{stable}$
BA .258 .030 0.019 0.016 732 AB
OBP .328 .025 0.019 0.022 442 PA
BB% 0.086 0.026 0.011 0.024 141 PA
K% 0.210 0.052 0.017 0.050 67 PA
H% 0.230 0.024 0.017 0.016 677 PA
HR% 0.034 0.013 0.007 0.011 291 PA
3.2.2 Confidence Intervals
Another way to understand statistic stabilization is through the lens of sequentially valid confidence intervals for player metrics. One straightforward approach is to construct these intervals using Hoeffding’s inequality. Since the outcome of each event (e.g., plate appearance) is modeled as $X_{i}\sim\text{Bernoulli}(p)$, we know that $0\leq X_{i}\leq 1$. By Hoeffding’s inequality, for any $v>0$, we have:
$\mathbb{P}(|\bar{X}_{t}-p|\geq t)\leq 2\exp(-2tv^{2}),$
where $\bar{X}_{t}$ is the sample mean after $t$ observations.
Setting $v=\sqrt{\frac{1}{2t}\log(2/\alpha)}$ for a desired confidence level $1-\alpha$, we obtain:
$\mathbb{P}\left(|\bar{X}_{t}-p|\geq\sqrt{\frac{1}{2t}\log(2/\alpha)}\right)\leq\alpha.$
Thus, a sequentially valid $(1-\alpha)$ confidence interval at time $t$ is given by:
$\left[\hat{p}-\sqrt{\frac{1}{2t}\log(2/\alpha)},\ \hat{p}+\sqrt{\frac{1}{2t}\log(2/\alpha)}\right].$
This framework connects naturally to our definition of stabilization: we can interpret a metric as stabilized once this confidence interval becomes narrower than a predefined tolerance. In contrast to RMSE-based thresholds, this approach provides probabilistic guarantees that remain valid at every point in time, making it well-suited for in-season monitoring of player metrics. Figure 1 plots Aaron Judge’s OBP after each game in the 2024 season, along with sequentially valid confidence intervals computed using Hoeffding’s inequality.
Figure 1: Sequentially valid confidence sequence for Aaron Judge’s OBP in the 2024 season. Each point corresponds to his OBP after a game (with the number of cumulative PAs on the x-axis). The error bars correspond to the 95% confidence interval constructed using Hoeffding’s inequality.
3.3 Extensions to other bounded metrics
While the examples above focus on binary outcomes modeled with Bernoulli distributions, many other baseball metrics are similarly bounded and can be treated using analogous methods. For instance, metrics such as average fastball velocity, launch angle, and exit velocity are continuous-valued but fall within well-defined ranges. In these cases, we can still apply concentration inequalities like Hoeffding’s by leveraging the fact that the data are bounded between known minimum and maximum values. This allows us to derive sequential confidence intervals and stabilization thresholds without making strong distributional assumptions.
For example, suppose we know that $X_{i}\in[a,b]$, then applying Hoeffding’s inequality yields:
$\mathbb{P}(|\bar{X_{t}}-\mu|\geq v)\leq 2exp(-\frac{2tv^{2}}{(b-a)^{2}})$
If we set $v=\sqrt{\frac{(b-a)^{2}}{2t}\log(2/\alpha)}$ we get that $\mathbb{P}(|\bar{X_{t}}-\mu|\geq v)\leq\alpha$. Thus, a sequentially valid confidence interval at time $t$ is $\bar{X_{t}}-\sqrt{\frac{(b-a)^{2}}{2t}\log(2/\alpha)},\bar{X_{t}}+\sqrt{\frac{(b-a)^{2}}{2t}\log(2/\alpha)}]$.
While Hoeffding-based intervals are a natural choice for bounded metrics, other valid approaches exist. For example, one can construct sequential confidence intervals using betting martingales (Waudby-Smith & Ramdas 2024), which offer nonparametric guarantees and are tightly connected to recent advances in online inference. Alternatively, one might use Gaussian intervals with corrections for multiplicity to maintain valid coverage over time.
4 Changepoint detection
In this section we present a tractable changepoint detection algorithm for player performance data. Changepoint detection aims to identify the time point, if one exists, at which the statistical distribution generating the data changes. Suppose we have a sequence of data $\{y_{1},...,y_{n}\}$. The goal is to identify a point $t$, if it exists, such that
$y_{1},...,y_{t}\sim\mathcal{P}\text{ and }y_{t+1},...,y_{n}\sim\mathcal{Q},\hskip 5.69054pt\mathcal{P}\neq\mathcal{Q}.$
Here, we are primarily concerned with the problem of identifying changes in the true mean of a given player’s performance metric. For example, detecting when a batter’s out-of-zone swing percentage decreases or when a pitcher’s fastball velocity increases.
4.1 Likelihood-Ratio Based Tests
A common method for detecting changepoints uses a likelihood-based framework (Hinkley 1970, Hinkley & Hinkley 1970). Consider a time-ordered sequence of data $y_{1:n}=(y_{1},\dots,y_{n})$. Let $p(y\mid\theta)$ denote the probability density function of $y$ given parameter $\theta$, and let $\hat{\theta}$ denote the maximum likelihood estimate for the full dataset under the null hypothesis of no changepoint.
Under the alternative hypothesis, there exists a changepoint $t$ such that $y_{1:t}$ follow $p(y\mid\theta_{0})$ and $y_{t+1:n}$ follow $p(y\mid\theta_{1})$. Let $\hat{\theta}_{0}^{(t)}$ and $\hat{\theta}_{1}^{(t)}$ be the MLEs fit to each respective segment. Then the log-likelihood under the alternative is:
$\ell(t)=\log(p(y_{1:t}\mid\hat{\theta}_{0}^{(t)}))+\log(p(y_{t+1:n}\mid\hat{\theta}_{1}^{(t)}))$
Under the null, the log-likelihood is
$).\log(p(y_{1:n}\mid\hat{\theta}$)).
Define
$).\lambda_{t}=\ell(t)-log(p(y_{1:n}\mid\hat{\theta}$)).
and take the test statistic as
$\lambda=\max_{t}\lambda_{t}.$
with the candidate changepoint
$t^{\prime}=\operatorname*{arg\,max}_{t}\lambda_{t}.$
If $\lambda>c$ for some threshold $c$ (chosen to control the Type I error rate), we declare $t^{\prime}$ to be a changepoint.
Choosing an appropriate value of $c$ can be challenging in practice (Shang et al. 2017). This issue frequently arises in sports applications, as common changepoint detection algorithms often overflag changepoints, leading to high false positive (Type I error) rates (see Sections 5 and 6 for examples). In a player performance context, some false positives may be acceptable to maintain power to detect true changes, since flagged changes would typically be reviewed manually by teams. However, excessive false positives can overwhelm limited time and resources for review. In practice, generic algorithms often produce so many changepoints that manual vetting becomes infeasible.
4.2 Split sample inference
To appropriately control the type I error rate while still maintaining computational tractability, we propose combining changepoint detection with split-sample inference (Cox 1975). Split-sample inference is a statistical technique where data is divided into two parts: one part to train on (i.e., detect changepoints) and the other part to test (i.e., test whether the average difference in the metric before and after the changepoint is significant). In our setting, we split the data into even and odd indices to ensure comparable sample sizes and reduce dependence between the training and testing sets. The odd-indexed observations will be used to calculate the potential changepoint location $t^{\prime}$, and the even-indexed observations will be used to test the null hypothesis $\theta_{0}=\theta_{1}$ (or, as described in Section 6, a shifted null $\theta_{1}=\theta_{0}+\Delta$ for a pre-specified $\Delta$). The single changepoint detection algorithm, incorporating likelihood-ratio based test statistics and split sample inference, is presented in Algorithm 1.
Algorithm 1 Single Changepoint Detection Across Players
1:Input: NN (number of players), $\alpha$ (significance level), and the statistic of interest.
2:for each player $i=1,\ldots,N$ do
3:  Obtain the time-ordered statistic sequence for player ii.
4:  Split the sequence into two disjoint sets: odd-indexed and even-indexed timepoints.
5:  Let $T_{odd}$ be the set of odd-indexed timepoints.
6:  for each $t\in T_{odd}$ do
7:   Compute the log-likelihood test statistic $\lambda_{t}$ (e.g., Equation 2 for binary data and Equation 6 for normally distributed data).
8:  end for
9:  Identify the candidate changepoint: $t^{*}=\arg\max_{t}\lambda_{t}$.
10:  Perform a hypothesis test at level $\alpha$ using the even-indexed timepoints.
11:  if the test is significant then
12:   Declare $t^{*}$ as the changepoint.
13:  else
14:   Declare no changepoint for player ii.
15:  end if
16:end for
4.3 Multiple changepoint detection
Algorithm 1 details our method for detecting a single changepoint, which is a common use case in sports settings. In practice, teams often want to identify a changepoint, flag it, and continue monitoring for the next changepoint as new data arrive. Algorithm 1 naturally extends to multiple changepoint detection via binary segmentation, which is helpful when scanning a long historical window for several performance shifts.
Binary segmentation works by repeatedly applying the single changepoint detection algorithm to successively smaller segments. Starting with the full series, we (1) run Algorithm 1 on the current segment; (2) if a significant changepoint is found, we split the segment at that location; and (3) recursively analyze the left and right subsegments. We stop splitting a segment when no significant changepoint is detected (at level $\alpha$) or when either child would be shorter than a minimum length $m$, which ensures the test in Algorithm 1 has adequate data on both sides of a candidate split. See Algorithm 2.
Because multiple tests are performed across segments, users may optionally control the overall false discovery rate (e.g., by adjusting $\alpha$ adaptively or applying a correction such as Bonferroni across detected splits).
In Section 5 and 6 we apply Algorithm 2 to detect changes in batter plate discipline and pitcher fastball velocity respectively. We chose these metrics because they are more directly under the batter’s or pitcher’s control and less influenced by the opposing team or ballpark. Code to implement all changepoint detection algorithms is available at https://github.com/akglazer/sports-changepoint.
Algorithm 2 Multiple Changepoint Detection Across Players (Binary Segmentation)
1:Input: NN (number of players), $\alpha$ (significance level), $m$ (minimum segment length), and the statistic of interest.
2:for each player $i=1,\ldots,N$ do
3:  Obtain the time-ordered statistic sequence $x_{i}$ of length $n_{i}$ for player ii.
4:  Initialize the set of detected changepoints $\mathcal{C}_{i}\leftarrow\varnothing$.
5:  Initialize a segment list $\mathcal{S}\leftarrow\{(1,n_{i})\}$.
6:  while $\mathcal{S}$ is not empty do
7:   Select a segment $(s,e)$ from $\mathcal{S}$.
8:   if es+1<2me-s+1<2m then
9:     continue $\triangleright$ not enough length to split into two parts of size $\geq m$
10:   end if
11:   Apply Algorithm 1 to the subsequence $x_{i}[s:e]$ to obtain a changepoint $\tilde{t}$ (local index in $[s,e]$).
12:   if no changepoint is returned at level $\alpha$ then
13:     Do not split $(s,e)$ further.
14:   else
15:     Convert to global index: $t^{*}\leftarrow s+\tilde{t}-1$.
16:     if $t^{*}-s+1\geq m$ then
17:      Add $(s,t^{*})$ to $\mathcal{S}$.
18:     end if
19:     if $e-t^{*}\geq m$ then
20:      Add $(t^{*}+1,e)$ to $\mathcal{S}$.
21:     end if
22:     $\mathcal{C}_{i}\leftarrow\mathcal{C}_{i}\cup\{t^{*}\}$.
23:   end if
24:  end while
25:  Output $\mathcal{C}_{i}$ for player ii.
26:end for
5 Detecting changes in batter plate discipline metrics
In this section we apply Algorithm 2 to detect changes in two batter plate discipline metrics: chase rate and whiff rate. First, consider chase rate. For a particular batter, let $\{y_{i}\}_{i=1}^{n}$ be the set of all $n$ pitches outside of the strike zone with
[…]
We want to determine whether there exists a timepoint $t^{\prime}$ such that
$\{y_{i}\}_{i=1}^{t^{\prime}}\sim\text{Bernoulli}(p_{1})\text{ and }\{y_{i}\}_{i=t^{\prime}+1}^{n}\sim\text{Bernoulli}(p_{2})\text{ where }p_{1}\neq p_{2}.$
We define a batter’s chase rate across ba+1b-a+1 timepoints spanning the $a$th (out of zone) pitch to the $b$th pitch to be:
$\bar{y}_{a:b}:=\frac{1}{b-a+1}\sum_{i=a}^{b}y_{i}$
Following the approach outlined in Section 4, for a timepoint $t$, we define
$\ell(t)=\log(\bar{y}_{1:t}^{\sum_{i=1}^{t}y_{i}}((1-\bar{y}_{1:t})^{\sum_{i=1}^{t}(1-y_{i})})+\log(\bar{y}_{t+1:T}^{\sum_{i=t+1}^{T}y_{i}}((1-\bar{y}_{t+1:T})^{\sum_{i=t+1}^{T}(1-y_{i})})$ (1)
and
$\lambda_{t}=\ell(t)-\log(\bar{y}_{1:T}^{\sum_{i=1}^{T}y_{i}}((1-\bar{y}_{1:T})^{\sum_{i=1}^{T}(1-y_{i})}).$ (2)
The potential changepoint is
$t^{\prime}=\operatorname*{arg\,max}_{t}\lambda_{t}.$
As in algorithm 1, we use split-sample inference to determine whether $t^{\prime}$ is a changepoint. In particular, we split $\{y_{i}\}_{i=1}^{T}$ into odd and even pitches: $y^{odd}=\{y_{i}\}_{i\text{ mod }2=1}$ and $y^{even}=\{y_{i}\}_{i\text{ mod }2=0}$. We use the odd pitches, $y^{odd}$, to calculate our potential changepoint $t^{\prime}$. Then, we use the even pitches $y^{even}$ to test the null hypothesis that the average chase rate after $t^{\prime}$ is the same as the average chase rate before $t^{\prime}$: $p_{1}=p_{2}$. Since the data are generated from a Bernoulli distribution, in the hypothesis testing step, we apply Fisher’s Exact Test.
We also apply the same procedure to batter’s whiff rate. For whiff rate, let $\{y_{i}\}_{i=1}^{n}$ be the set of all $n$ pitches that a particular batter swings at. Then:
[…]
Using this definition of $y_{i}$, we apply the same procedure as previously described for chase rate.
5.1 Plate discipline results
Using data from the 2023 and 2024 MLB seasons, we ran Algorithm 2 on all batters who faced at least 100 out-of-zone pitches and who had at least 100 swings for chase rate and whiff rate respectively. Table 2 summarizes the results. Sample splitting reduces the number of batters flagged for changepoints by 89% for chase rate and 88% for whiff rate.
Table 2: Number of batters flagged for changepoints in chase rate and whiff rate using Fisher’s Exact Test with $\alpha=0.05$, with and without sample splitting. In the no sample splitting condition, Fisher’s Exact Test is applied to the same data used to detect potential changepoints.
Metric Batters Considered Changepoints (Sample Splitting) Changepoints (No Sample Splitting)
Chase Rate 647 50 (8%) 465 (72%)
Whiff Rate 687 58 (8%) 479 (70%)
The MLB regular season ran from March 30 to October 1 in 2023 and March 20 (for a two game series in South Korea between the Padres and Dodgers before the rest of the games began on March 28) to September 30 in 2024. For chase rate and whiff rate respectively, 62% and 64% of the flagged changepoints occured in May through August.
As an example, Figure 2 plots a rolling average of Brice Turang’s swinging strike rate for the 2023 season. The flagged changepoint is marked by the vertical dashed red line. While the changepoint detection algorithm operates directly on the binary pitch-level data, the rolling average is shown here as a useful way to visualize trends in a rate statistic. Brice Turang saw an improvement to his whiff rate from the 2023 to 2024 season, with his whiff rate decreasing from 21.7% to 13.8%. Using our changepoint detection algorithm, we are able to flag this change at the end of July 2023.
Figure 2: Brice Turang’s rolling average whiff rate, calculated over the last 50 swings, across the 2023 MLB season. The vertical red dashed line denotes the detected changepoint.
6 Detecting changes in pitcher fastball velocity
Next we consider the problem of detecting changes in a pitcher’s primary fastball velocity. We assume that the velocity of a pitcher’s $i$th fastball follows a normal distribution, consistent with empirical pitch-level data (MLB Advanced Media 2025): $y_{i}\sim\mathcal{N}(\mu,\sigma^{2})$ for some mean $\mu$ and variance $\sigma^{2}$.
Let
$\hat{\mu}_{a:b}:=\frac{1}{b-a+1}\sum_{i=a}^{b}y_{i}$
and
$\hat{\sigma}^{2}_{a:b}:=\frac{1}{b-a+1}\sum_{i=a}^{b}(y_{i}-\hat{\mu}_{a:b})^{2}.$
Then, for a timepoint $t$, we can define
$\ell(t)$ $=[-\frac{t}{2}\log(2\pi\hat{\sigma}^{2}_{1:t})-\frac{t}{2}]+[-\frac{T-t}{2}\log(2\pi\hat{\sigma}^{2}_{t+1:T})-\frac{T-t}{2}]$
$=-\frac{t}{2}\log(2\pi\hat{\sigma}^{2}_{1:t})-\frac{T-t}{2}\log(2\pi\hat{\sigma}^{2}_{t+1:T})-\frac{T}{2}$ (3)
and
$\lambda_{t}$ $=\ell(t)-[-\frac{T}{2}\log(2\pi\hat{\sigma}^{2}_{1:T})-\frac{T}{2}]$
$=-\frac{t}{2}\log(2\pi\hat{\sigma}^{2}_{1:t})-\frac{T-t}{2}\log(2\pi\hat{\sigma}^{2}_{t+1:T})+\frac{T}{2}\log(2\pi\hat{\sigma}^{2}_{1:T}).$ (4)
The potential changepoint is $t^{\prime}=\operatorname*{arg\,max}_{t}\lambda_{t}$.
Again, we implement split sample inference according to Algorithm 1, with one important modification. Here, we test for a changepoint under a null hypothesis that allows for a location shift in the parameter of interest. Let $\theta_{0}$ and $\theta_{1}$ denote the average metric values before and after the candidate changepoint, respectively. Under the null hypothesis, we assume
$\theta_{1}=\theta_{0}+\Delta,$
where $\Delta$ is a pre-specified shift. We implement a two-sample permutation test that incorporates this shift directly. This approach allows for flexible hypothesis testing, e.g., one can choose to flag only changepoints where the post-change fastball velocity differs significantly by at least 1 mph (by setting $\Delta=1$).
For metrics such as fastball velocity, where within-pitcher variability is low, even small fluctuations can appear statistically significant, especially with large sample sizes. However, these minor changes are often not meaningful from a baseball perspective. By incorporating a location shift $\Delta$ into the null hypothesis (e.g., testing whether the post-change average differs by more than 1 mph), we focus detection on changes that are practically significant. This approach reduces false positives and ensures that only meaningful deviations, such as those indicative of injury, mechanical adjustment, or performance decline, are flagged as changepoints. For low-variance metrics like velocity, where the difference between statistical and practical significance is especially pronounced, this modification improves interpretability and aligns inference with domain expertise.
6.1 Fastball velocity results
We apply Algorithm 2 using a two-sample permutation test with $\alpha=0.05$ to detect changes in four seam fastball velocity for starting pitchers in the 2024 MLB season. We define a starter as any pitcher that pitched more than 4 innings in an outing more than 10 times and entered in the 5th inning or later fewer than 5 times. We look at all pitchers with at least 100 four-seam fastball pitches.
Table 3 summarizes the results. While the impact of sample splitting is not as drastic as with the plate discipline metrics considered in the previous section, sample splitting still reduces the number of flagged changepoints. Most notably, running the permutation test with shift parameter equal to one substantially reduces the number of detected changepoints.
Table 3: Number of starters ($N=143$) in the 2024 MLB season flagged for changepoints in four-seam fastball velocity applying Algorithm 2 using a two-sample permutation test with $\alpha=0.05$, with and without sample splitting. In the no sample splitting condition, the permutation test is applied to the same data used to detect potential changepoints.
Threshold Sample Splitting Changepoints
0 Yes 130 (91%)
0 No 142 (99%)
1 Yes 22 (15%)
1 No 30 (21%)
Next, we consider starters across the 2023 and 2024 MLB seasons and run Algorithm 2 with $\alpha=0.05$ and two-sample permutation test with shift parameter equal to one. Of the 174 starters considered, we flag 49 changepoints across 40 starters of which 21 (43%) occur between September 1, 2023 and May 1, 2024 (i.e., during or near the offseason).
7 Ground-truth evaluation: pitcher role changes
One difficult aspect of changepoint detection in sports, and otherwise, is that it can be challenging or even impossible to acquire ground-truth data, because often we do not know if a change actually did occur. In baseball, true changes in performance can occur for a multitude of reasons. For example, a pitcher’s fastball velocity may
  • increase as he builds strength
  • increase due to a biomechanical change he made in his approach
  • decrease as he ages
  • decrease due to an injury
  • increase or decrease due to a change in role
It is difficult to construct a ground truth changepoint dataset for many of these types of changes because we do not know, for example, that a pitcher’s fastball velocity will definitely decrease as he ages from 32 to 33. We often also do not know if a pitcher has made a biomechanical change (unless, for example, they tell the media or we have inside information). However, the last bullet point provides an avenue for constructing a ground-truth dataset to test our changepoint detection algorithm, as we know when a pitcher experiences a role change (e.g., from starter to reliever). In this section we evaluate our algorithm’s ability to detect known changes in pitcher statistics due to a role change using a curated ground-truth set of pitcher role changes.
We construct a dataset of pitchers, in the 2023 and 2024 MLB seasons, that switched from a relief to starting pitcher role to tune and test our changepoint detection algorithms. We consider fastball velocity, as it is well established that pitchers tend to lose velocity when transitioning from a relief role to a starting role (and gain velocity when moving from starter to reliever), since outings with a higher pitch volume typically require pitchers to reduce their intensity in order to remain effective over a longer duration compared to shorter relief appearances (Choi 2022, Palattella 2023).
Pitchers switch roles for a variety of reasons. For example, they might switch from pitching as a starter to as a reliever due to reduced stamina, struggles after the first or second time through the batting order, repeated injuries, roster construction issues, or significantly better performance in shorter outings. On the other hand, pitchers might switch from pitching as a reliever to as a starter due to development of another pitch, success in longer relief stints, a mechanical adjustment, or a team need for more starters.
We focus on building a groundtruth dataset around reliever to starter transitions as relievers frequently transition to a starting role due to improved performance and stamina, but face a drop in velocity after transitioning due to the endurance required to pitch through longer outings. On the other hand starters transitioning to relievers may switch roles due to a decline in performance and the boost to their velocity that comes from the role change might be washed out due to their performance decline.
To identify players that transition from a relief to starting role, we look at players that entered the game in the 5th inning or later (indicative of relief outings) more than 10 times and pitched more than 4 innings (indicative of a start) more than 10 times across the 2023 and 2024 seasons. We then filtered this list to only consider pitchers that increased their average number of innings pitched from 2023 to 2024 (i.e., moved from a reliever to starter). We also excluded pitchers that transitioned roles more than once during this time period. We confirmed and supplemented this list in consultation with news articles highlighting reliever to starter transitions (Murphy 2024, Adams 2024).
The players listed in Table 4 were included in this “ground-truth” dataset. The changepoint algorithm searched for changes in the velocity of their primary fastball (sinker for Jordan Hicks, Michael King and Jose Soriano; four-seam for the other pitchers) in the 2023 and 2024 seasons. We implemented Algorithm 2 with $\alpha=0.05$ and increased the shift parameter of the permutation test until we were no longer able to detect fastball velocity changes for any player. We were able to successfully flag the velocity changepoint for 91% (10/11) of the pitchers. For 64% (7/11) of the starter to reliever transitions, a fastball velocity change was flagged with shift parameter equal to 1 ($\Delta=1$). Results are displayed in Table 4.
Table 4: Pitchers that transitioned from a relief to starting role at some point in the 2023 or 2024 MLB season. The “Changepoint flagged?” column indicates whether a changepoint was flagged in their exit velocity for their primary fastball (sinker for Jordan Hicks, Michael King and Jose Soriano; four-seam for all other players listed), and “Max CP Threshold” gives the maximum shift parameter ($\Delta$) that still flags a changepoint. Across all values of the shift parameter, Algorithm 2 was implemented with $\alpha=0.05$ and two-sample permutation tests.
Pitcher name MLB ID Changepoint flagged? Max CP Threshold
Jordan Hicks 663855 Yes 5
Reynaldo Lopez 625643 Yes 2
Ronel Blanco 669854 Yes 0.5
Michael King 650633 Yes 1
Zack Littell 641793 Yes 1
Jose Soriano 667755 No NA
Garrett Crochet 676979 Yes 0.5
Tyler Alexander 641302 Yes 0.5
Andre Pallante 669467 Yes 1
Cole Ragans 666142 Yes 1
Sean Manaea 640455 Yes 1
The largest change in fastball velocity was from pitcher Jordan Hicks. The only pitcher where a change in fastball velocity was not detected was Jose Soriano. Figure 3 plots fastball velocity by pitch for both Hicks and Soriano. Horizontal blue lines represent the mean sinker velocity before and after their role transition. While the average difference is quite apparent for Hicks (5.5 mph difference), it is negligible for Soriano (0.12 mph difference) making it clear why the algorithm did not flag a change in sinker velocity for Soriano.
Figure 3: Jordan Hick’s and Jose Soriano’s sinker velocity by pitch for MLB games in 2023 and 2024. Each plot features two horizontal blue lines. The leftmost and rightmost horizontal lines represent the average sinker velocity for each pitcher as a reliever and starter respectively. Hick’s average sinker velocity decreases by 5.5 mph after the transition. Soriano’s average sinker velocity increases by 0.12 mph after the transition.
8 Discussion
In this paper we present methods for quantifying metric stabilization, as well as tractable algorithms for detecting changepoints in binary and continuous player performance metrics. In particular, we propose the use of a likelihood based changepoint detection algorithm combined with split sample inference, using either nonparametric tests or tests appropriate to the underlying data distribution. Furthermore, hypothesis tests can incorporate a shift parameter, allowing users to only flag changes larger than a pre-specified magnitude. We demonstrate the application of this algorithm on batter plate discipline metrics and pitcher fastball velocity. Our method is able to flag velocity changes in 91% of pitchers that transitioned from a relief to starting role in the 2023 or 2024 MLB season. Furthermore, we find that for some metrics more than 60% of detected changes occur in-season.
While this work draws on existing changepoint detection literature, methods are combined in a way specifically tailored to this setting. In particular, the inclusion of split-sample inference with hypothesis tests that allow for a shift parameter allows users to better control the Type I error rate as well as have more control over the magnitude of changes detected. Furthermore, existing R packages for changepoint detection generally do not handle binary data well, as calculations are typically based on the normal distribution (Killick & Eckley 2014, Lindeløv 2024). We provide an R implementation of all methods at https://github.com/akglazer/sports-changepoint.
We chose to focus on plate discipline metrics and fastball velocity as those metrics are less likely to be influenced by other factors such as the opponent and ballpark. However, fastball velocity can experience small fluctuations throughout a game as a pitcher fatigues. Future research could investigate the impact of fatigue on pitcher performance and integrate this into changepoint detection methods.
In this paper we primarily focus on detecting changes in means, however, the methods could easily be extended to look for changes in variance or other distributional characteristics. Furthermore, a nonparametric approach could be employed instead of the likelihood ratio-based method when the underlying data distribution is unknown. For example, one could use a divergence measure, as in Matteson & James (2014), instead of a likelihood ratio test statistic.
While we focus on split sample inference as a way to reduce the false positive rate, alternative approaches could also be considered. One such method, proposed by Jewell et al. (2022), evaluates the probability, under the null hypothesis of no true change in mean at a candidate changepoint, that the observed difference in means between segments on either side of a candidate changepoint could arise by chance. Specifically, it conditions on the set of changepoints and assesses whether the observed change is unusually large relative to all datasets that would yield the same changepoint configuration.
We could adapt this method to the binary setting by generating datasets $y^{\prime}=\{y^{\prime}_{i}\}_{i=1}^{T}$ such that $y^{\prime}_{i}\sim\text{Bernoulli}(\bar{y}_{1:T})$ and restricting attention to those in which a changepoint $\lambda^{\prime}$ falls within a window around the observed location, e.g., $\lambda^{\prime}\in[\lambda-50,\lambda+50]$. We could then calculate $p=Pr\{\bar{y^{\prime}}_{\lambda^{\prime}+1:T}-\bar{y^{\prime}}_{1:\lambda^{\prime}}\geq\bar{y}_{\lambda+1:T}-\bar{y}_{1:\lambda}\}$, declaring $\lambda$ a changepoint if $p\leq\alpha$, where $\alpha$ is the desired significance level. A major challenge of this method, particularly at scale, is its computational cost, making it difficult to apply in real-time settings where analysts may wish to monitor many metrics across thousands of players daily. Nonetheless, future research could explore trade-offs between power and computation time, as well as potential algorithmic speed-ups or approximations that make this approach more feasible in practice.
While we focus on baseball applications in this paper, this work could easily be applied to other areas where it is of interest to detect changes in performance, e.g., manufacturing (Shang et al. 2017), healthcare or finance.